OpenStudy (anonymous):

Solve the system of equations -2x - 2y = -4 2x - y = -5

OpenStudy (anonymous):

A. (-1,3) B. (3,-1) C. (-1,-3) D. (-3,-1)

OpenStudy (texaschic101):

add them... -2x - 2y = -4 2x - y = -5 ------------ First, what does -2x + 2x equaal ?

OpenStudy (texaschic101):

*equal -- typo

OpenStudy (anonymous):

0?

OpenStudy (texaschic101):

correct. Now what does -2y + (-y) equal ?

OpenStudy (anonymous):

I'm sorry, I have no clue.

OpenStudy (texaschic101):

when the signs are the same, negative or positive, you add them. -2y + (-y) =- 3y. Now if the signs would have been different, and the problem would have been 2y + (-y), then you would subtract and take the sign of the largest number and this one would equal 1y or just y. Anyways, the signs were the same, both negative, so you just add them. Do you understand ?

OpenStudy (anonymous):

Yes, thank you.

OpenStudy (texaschic101):

now , lets finish... -4 + (-5).....what does that equal ?

OpenStudy (anonymous):

-9?

OpenStudy (texaschic101):

correct. -2x - 2y = - 4 2x - y = - 5 -------------(add) 0 _ 3y = - 9 -3y = - 9 Do you know how to solve for y ?

OpenStudy (anonymous):

Um..*sigh* Sort of. I suck at math.

OpenStudy (texaschic101):

-3y = - 9 we need to get y by itself. To do this, we do the opposite of what the problem is doing. Let me try to show you...um.....-3y means -3*y. It is multiplying. We do the opposite, and divide. If we divide by -3 (on both sides), it will get y by itself. -3y/-3 = -9/-3. the -3's divide to make 1y = -9/-3 (keep in mine, when you are dividing by two numbers with the same sign, it changes the sign to a positive. So, -9/-3 = 3. Can you kinda understand that ? I am not real good at explaining. Anyways....y = 3

OpenStudy (anonymous):

So, it can't be C or D..Right?

OpenStudy (texaschic101):

right. we know that y = 3. To find x, we sub 3 in for y in either of the original equations. 2x - y = - 5 2x - 3 = - 5 (now you add 3 to both sides to get 2x by itself) 2x - 3 + 3 = - 5 + 3 (combine like terms) 2x = - 2 (divide by 2 to get x by itself) 2x/2 = - 2/2 x = - 1 Can you see what I did ? Does it make sense ?

OpenStudy (anonymous):

Okay. I get it. Simple equations! Thank you so much! Would it be B then?

OpenStudy (texaschic101):

lets see.....x = -1 and y = 3. In the ordered pair (x,y).....x comes first.So the answer is (-1,3)....which is A. Remember, x comes first. You will need to study more on your signs, but other then that, you got this. Good job.

OpenStudy (texaschic101):

Do you have any more questions ?

OpenStudy (texaschic101):

well.....um...if so, tag me and I might can help you. If not, have a nice day :)

OpenStudy (anonymous):

Dang, I remember that now. (x,y) Thank you so much!!!!! @texaschic101